OpenStudy (anonymous):

Assume that 2 digits are selected at random from the set {1,3,5,7,9} and are arranged in random order. What is the probability that the resulting 2-digit number is less than 90?

OpenStudy (anonymous):

from 2 digits are selected, each number can repeat or no?

OpenStudy (anonymous):

i'm going to assume they can repeat.

OpenStudy (anonymous):

OK..

OpenStudy (anonymous):

for solution : Probability of an event happening = Number of ways it can happen/Total number of outcomes = 4(5)/5(5) = 20/25 = 4/5